### v2

# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

Abstract
In this paper, we propose a novel convex optimization framework to solve the optimal covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) with chance constraints. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we use Gaussian approximations to incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under Gaussian noise and Markovian jump dynamics.
I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9]. Such applications typically require path constraints due to safety considerations and physical limitations and have previously been incorporated in the form of probabilistic chance constraints[10]. [11] proposed a lossless convex relaxation of the chance-constrained covariance steering problem, which was later shown [12] to outperform existing methods [4, 6] in computation time.
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a class of hybrid systems that involve state-triggered switching. However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS). MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16]. [17] is an excellent reference on the classical results for discrete-time MJLS.
Constrained control of MJLS has been an active area of research, e.g. with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20]. Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic path planning approaches which face abrupt changes in dynamics, such as spacecrafts[22] and robots[20] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
Notation
[…]
II MJLS Dynamics
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
$x_{k+1}=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (1)
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in $\Omega=\{1\ldots N\}$ with transition probability matrix $\mathbb{P}$ whose ($i$,$j$)th entry is $p_{ij}$. We define the following variables to study MJLS dynamics:
$\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\mathcal{X}_{k}(i)\triangleq\mathbb{E}\left[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}\right],$
$X_{k}\triangleq\mathbb{E}[x_{k}x_{k}^{\top}]=\sum_{i\in\Omega}\mathcal{X}_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
We assume that the initial probability distribution $\rho_{0}\in\mathbb{R}^{1\times N}$ is given. Therefore, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
Remark 1.
Note that $x_{k}$ is independent of $r_{k}$. Therefore, $q_{k}(i)=\mathbb{E}[x_{k}]\mathbb{E}[\mathds{1}_{r_{k}=i}]=\mu_{k}\rho_{k}(i)$, and $\mathcal{X}_{k}(i)=\rho_{k}(i)X_{k}$.
We assume a control policy of the following form:
$u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k})$ (2)
Assumption 1.
$\rho_{k}(i)>0,\;\forall k\in\mathbb{Z}_{0}^{T},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
Assumption 2.
The covariance of the given MJLS is positive definite at all time steps, i.e. $\Sigma_{k}\succ 0$ $\forall k\in\mathbb{Z}_{0}^{T}$.
II-A Mean Propagation
We now focus on MJLS mean dynamics. We know:
$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
$=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
$\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
$=\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (3)
Using the definition of $\mu_{k}$, eq. 3 can also be written as:
$\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (4)
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy eq. 4.
II-B Covariance Propagation
For covariance propagation, we first consider $\mathcal{X}_{k+1}(j)$:
$\mathcal{X}_{k+1}(j)=\mathbb{E}\left[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}\right]$ (5)
$=\sum_{i\in\Omega}\left\{p_{ij}\mathbb{E}\left[\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)\right.\right.$
$\left.\left.\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)^{\top}\mathds{1}_{r_{k}=i}\right]\right\}$
Using identities derived in the appendix, we can simplify eq. 5, and substitute that in $\Sigma_{k+1}=\sum_{j\in\Omega}\mathcal{X}_{k+1}(j)-\mu_{k+1}\mu_{k+1}^{\top}$ to get:
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
$\left.+A_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K_{k}^{\top}(i))B^{\top}_{k}(i)+A_{k}(i)\mu_{k}c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K^{\top}_{k}(i))^{\top}A_{k}^{\top}(i)+B_{k}(i)\overline{u}^{\top}_{k}(i)c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)\Sigma_{k}K_{k}^{\top}(i))B_{k}^{\top}(i)+c_{k}c_{k}^{\top}\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
Remark 2.
Separately analyzing some terms inside the summation in eq. 6:
$A_{K}(i)\mu_{k}\mu_{k}^{\top}A_{K}^{\top}(i)+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}$
$+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)$
$+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}B_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
$=(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})^{\top}$
Note these terms are linear in $p_{ij}$ while $\mu_{k+1}\mu_{k+1}^{\top}$ involves the product $p_{ij}p_{ij}$, explaining why they don’t cancel out. This demonstrates why the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
III Unconstrained MJLS Covariance Steering
III-A Optimal Control Problem
The unconstrained optimal control problem is given by:
$\min$ $\;J=\sum_{k=0}^{T-1}J_{k}$ (7)
$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},$
$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+G_{k}(r_{k})c_{k}$
$x_{0}$ $\sim\mathcal{N}(\overline{\mu}_{0},\overline{\Sigma}_{0})$
$\mathbb{E}[x_{T}]=$ $\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$
where $J_{k}=\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ with $Q_{k}\succeq 0$ and $R_{k}\succ 0$, $\forall k\in\mathbb{Z}_{0}^{T-1}$. Simplifying $J_{k}$, we get:
$J_{k}$ $=\sum_{i\in\Omega}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
$=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (8)
$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
III-B Convex Relaxation
The nonlinear terms in eq. 6 result in a non-convex optimization landscape for eq. 7. To solve it in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 6:
$L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (9)
Rearranging eq. 6 using the above-mentioned terms,
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)L^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B^{\top}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}\right.$
$\left.+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)]\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a finite sequence of $(2N+1)$-tuples $\{\Sigma_{k},L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
$\Sigma_{k+1}=F_{\tau}(\Xi)$ (11)
where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 9 as:
$C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (12)
which can be recast as a LMI using Schur’s complement:
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}\succeq 0$ (13)
We use these new variables in the objective function as follows:
$J$ $=J(\tau)+J(\Xi)$ (14)
where
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right]$ (15a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}({\Sigma}_{k}Q_{k}+Y_{k}(i)R_{k})\right]$ (15b)
III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 1 and then uses its solution to solve Problem 2.
Problem 1.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i)}J(\tau)$ (16a)
[…]
$\mu_{0}=\overline{\mu}_{0},\;\mu_{T}=\mu_{F}$ (16c)
Problem 2.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J(\Xi)$ (17a)
[…]
$H_{K}^{\Sigma}=\Sigma_{k+1}$ $-F_{\tau}(\Xi)=0$ (17b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (17c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (17d)
Assumption 3.
Slater’s condition is satisfied for Problem 2, i.e. there exists a strictly feasible solution for Problem 2.
Lemma 1 (Lemma 1 in [11]).
Let $A$ and $B$ be $n\times n$ symmetric matrices with $A\succeq 0,B\preceq 0$, and $\mathrm{Tr}(AB)=0$. If $B$ has at least one nonzero eigenvalue, then, $A$ is singular.
Theorem 1.
Under Assumptions 1, 2 and 3, the optimal solution to Problem 2 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
Proof.
For Problem 2, we can define the Lagrangian as:
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (18)
$+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is defined in eq. 12, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 17b which shares the same symmetry properties with eqs. 11 and III-B. $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (19a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (19b)
$\mathrm{Tr}($ $M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0$ (19c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 19c. Rearranging eq. 19a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 9, we simplify eq. 19b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
Remark 3.
If we attempt to carry out similar relaxations for the mean quadratic terms to relax the covariance propagation equation (6) without assuming the mean trajectory $\tau$ is given, we find that it is not possible to prove that these relaxations are lossless. In fact, we get a condition where $M_{k}^{(3)}=-M^{(4)}_{k}$, where $M_{k}^{(3)}$ and $M^{(4)}_{k}$ are Lagrange multipliers for respective relaxations. Since they are both positive semi-definite, it implies $M_{k}^{(3)}=M^{(4)}_{k}=0$. Thus, we cannot use our existing arguments to prove the relaxation is lossless. Numerical simulations also demonstrated this lossy behavior, motivating our two-step sequential mean-covariance optimization approach.
IV Chance-Constrained MJLS Covariance Steering
We now consider chance constraints for both state and control. For state, we consider intersection of hyperplane constraints and tube-like constraints around the mean trajectory. Mathematically, they’re expressed as:
$P(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (20a)
$P(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (20b)
Likewise, for control, we consider
$P(p^{\top}_{j}u_{k}(i)+q_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (21a)
$P(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (21b)
Previous works [23, 8] have focused on deterministic and convex reformulations of eqs. 20 and 21, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [24] between the MJLS distribution and the approximating Gaussian. It has been shown previously [25] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 20 and 21.
IV-A Deterministic Formulation
Lemma 2 (Lemma 2 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$ and $\epsilon\in(0,0.5)$. Then, $\mathbb{P}[a_{j}^{\top}{\xi}+b_{j}\leq 0,\;\forall j]\geq 1-\epsilon$ is implied by
$a_{j}^{\top}\bar{{\xi}}+b_{j}+m_{\mathcal{N}}(\epsilon_{j})\sqrt{a_{j}^{\top}P_{\xi}a_{j}}\leq 0,\;\forall j,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (22)
where $m_{\mathcal{N}}(\epsilon)=Q_{X\sim\mathcal{N}}(1-\epsilon)$ denotes the quantile function of the standard normal distribution at probability $1-\epsilon$.
Lemma 3 (Lemma 3 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$, $\epsilon\in(0,1)$, and $\gamma>0$. Then, $\mathbb{P}[\|{\xi}\|\leq\gamma]\geq 1-\epsilon$ is implied by
$\|\bar{{\xi}}\|+m_{\chi^{2}}(\epsilon,n_{\xi})\sqrt{\lambda_{\max}(P_{\xi})}\leq\gamma$ (23)
$m_{\chi^{2}}(\epsilon,n_{\xi})=\sqrt{Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)}$. Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
Note that while eqs. 22 and 23 are convex in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (24)
$m^{2}_{\chi^{2}}(\epsilon,n_{\xi}){\lambda_{\max}(P_{\xi})}\leq(\gamma-\|\bar{\xi}\|)^{2}$ (25)
which are convex in the covariance $P_{\xi}$.
IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [26, 27]. We now define the relaxed mean and covariance problems as follows:
Problem 3.
Given a covariance trajectory $\Xi$, solve:
$\min_{\mu_{k},\overline{u}_{k}(i),\beta_{m}}J_{\mu}^{c}$ (26a)
[…]
$a_{j}^{\top}\mu_{k}$ $+b_{j}+m_{\mathcal{N}}(\delta_{x,j})\sqrt{a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k}$ (26c)
$p_{j}^{\top}\overline{u}_{k}(i)$ $+q_{j}+m_{\mathcal{N}}(\delta_{u,j}(i))\sqrt{p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (26d)
$\|\overline{u}_{k}(i)\|+$ $m_{\chi^{2}}(\epsilon_{u}(i),n_{u})\sqrt{\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (26e)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq. 15a. $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight corresponding to the slack variable vector 1-norm $\|\beta_{m}\|_{1}$ in $J_{\mu}^{c}$.
Problem 4.
Given a mean trajectory $\tau$, solve:
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i),\zeta_{m}}J^{c}_{\Sigma}\hskip 56.9055pt$ (27a)
[…]
$H_{K}^{\Sigma}=\Sigma_{k+1}-F_{\tau}(\Xi)$ $=0$ (27b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (27c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (27d)
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (27e)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $\leq\zeta_{2,j},$ (27f)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (27g)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (27h)
where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq. 15b. $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$. $\alpha_{m}>0$ is the weight for slack variable vector 1-norm $\|\zeta_{m}\|_{1}$ in $J_{\Sigma}^{c}$.
Assumption 4.
Slater’s condition is satisfied in Problem 4 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 27c remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 4.
Theorem 2.
Under Assumptions 1, 2 and 4, the optimal solution to Problem 4 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
Proof.
We can define the Lagrangian for Problem 4 as follows:
$\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl{[}\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (28)
$+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right.$
$\left.+\lambda_{u,h}(i)e_{u,h}(i)\right\}\Bigr{]}+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
where $C_{k}(i)$ is given in eq. 12 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 27e, 27f, 27g and 27h respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (29a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (29b)
$\mathrm{Tr}(M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0,$
$\hskip 56.9055pt\lambda_{x,n},\lambda_{x,h},\lambda_{u,h},\lambda_{u,n}\geq 0$ (29c)
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 29c. Rearranging eq. 29a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 9, we simplify eq. 29b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the left hand side as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\Omega$. ∎
IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 1) and then solve Problem 4 and Problem 3 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
Algorithm 1 MJLS Chance-constrained covariance steering
1:Set tol, initial weights $\{\alpha_{m}\}$, and $\eta>1$;
2:Initialize mean trajectory $\tau$ by solving Problem 1;
3:while NotConverged do
4:     Solve Problem 4 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:     Compute new mean trajectory $\tau$ by solving Problem 3 using revised covariance trajectory $\Xi$ from previous step ;
6:     Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
7:     if $\max_{m}\|\beta_{m}\|_{\infty}$$>$tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ $>$tol then
8:         increase weights $\{\alpha_{m}\}$ by a factor $\eta$;
9:     else
10:         Solution converged. break;
11:return $\tau$ and $\Xi$
V Numerical Simulations
We use the following dynamics:
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}0.1&0\\ 0&0.1\end{bmatrix}$
$A_{k}(2)$ $=\begin{bmatrix}0.2&0.1\\ -0.5&0.1\end{bmatrix}$ $B_{k}(2)$ $=\begin{bmatrix}0&1\\ -1&2\end{bmatrix}$ $G_{k}(2)$ $=\begin{bmatrix}0.5&0\\ 0&0.5\end{bmatrix}$
$\mathbb{P}$ $=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$. We set the cost weight matrices $Q_{k}=0$ and $R_{k}=I$ $\forall k\in\mathbb{Z}_{0}^{T-1}$.
We use the CVXPY package[28] with MOSEK solver[29] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=1e2$, with $\eta=5$ and $tol=1e-6$. It takes 3.43 seconds and $3$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 1(a) and 1(b) plot 1000 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. Both figures use the control norm chance constraint. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints, enforced using the state half-space constraint. Figure 2 shows the control norm history for the Monte Carlo simulations. None of the samples disobeyed the state chance constraints, while the control norm exceeded $u_{\max}$ in $1.3\%$ of the total sample size, which meets the prescribed risk bound of $5\%$. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 1 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 1(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
(a) unconstrained case
(b) with chance constraints
Figure 1: State trajectories for Monte Carlo simulations with the state chance constraint $P(x_{1}\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 5)\leq 0.95$.
Figure 2: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{1}\geq-10)\geq 0.95$.
VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented.
Appendix A MJLS identities
We show some of the identities which are used in deriving the covariance propagation equation for MJLS.
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\rho_{k}(i)\left[\mu_{k}\overline{u}^{\top}_{k}(i)+{\Sigma}_{k}K^{\top}_{k}(i)\right]$ (30)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
$=\overline{u}_{k}(i)c_{k}^{\top}\mathbb{E}[\mathds{1}_{r_{k}=i}]+0=\overline{u}_{k}(i)c_{k}^{\top}\rho_{k}(i)$ (31)
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))(\overline{u}_{k}(i)+$
$\hskip 142.26378ptK_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}\mathds{1}_{r_{k}=i}]-$
$\mu^{\top}_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])K_{k}^{\top}(i)+K_{k}(i)(\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]-\mu_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])\overline{u}_{k}^{\top}(i)$
$\hskip 42.67912pt+K_{k}(i)\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (32)

### v3

@@
 I Introduction
-Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function.
-It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
+Covariance steering is an important problem of designing a control policy that steers a system from an initial distribution to a target distribution while minimizing a cost function.
+It was first proposed in [1] as an infinite horizon problem and has attracted significant interest recently[2, 3, 4, 5].
 The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9].
@@
 Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored.
-[13] studies covariance steering for a class of hybrid systems that involve state-triggered switching.
-However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS).
-MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain.
-They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16].
-[17] is an excellent reference on the classical results for discrete-time MJLS.
-Constrained control of MJLS has been an active area of research, e.g.
+[13] studies covariance steering for a subclass of hybrid systems but don’t extend to Markov jump linear systems (MJLS).
+MJLS is a type of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain.
+They have been extensively studied in the literature[14] with applications in fault-tolerant control [15], networks [16], and macroeconomic models [17].
+Constrained control of MJLS has been an active area of research, and has benefited from the enormous progress in constrained control of single mode systems [bemporad_explicit_2002, 18] e.g.
 with state and control polytopic constraints for linear time-invariant[19] and periodic systems[20], and with second-moment constraints[21].
 Chance constraints have previously been incorporated for MJLS[22], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons.
-Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner.
+Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints in a computationally tractable manner.
 To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
@@
 […]
-II MJLS Dynamics
+II Problem Statement
 We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
@@
 $\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
-$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\mathcal{X}_{k}(i)\triangleq\mathbb{E}\left[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}\right],$
-$X_{k}\triangleq\mathbb{E}[x_{k}x_{k}^{\top}]=\sum_{i\in\Omega}\mathcal{X}_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
-We assume that the initial probability distribution $\rho_{0}\in\mathbb{R}^{1\times N}$ is given. Therefore, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
-Remark 1.
-Note that $x_{k}$ is independent of $r_{k}$. Therefore, $q_{k}(i)=\mathbb{E}[x_{k}]\mathbb{E}[\mathds{1}_{r_{k}=i}]=\mu_{k}\rho_{k}(i)$, and $\mathcal{X}_{k}(i)=\rho_{k}(i)X_{k}$.
-We assume a control policy of the following form:
-$u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k})$ (2)
+$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
+The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$.
+For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
+We approximate $q_{k}(i)$ by $\rho_{k}(i)\mu_{k}$, which was found to be a reasonable approximation numerically.
+Similarly, we approximate $\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]$ by $\rho_{k}(i)\Sigma_{k}$.
+The appendix section elaborates on this approximation and the consequent justification.
+We now define the MJLS covariance steering optimal control problem below:
+Problem 1.
+Solve:
+$\min_{x_{k},u_{k}(i)}$ $\;J=\sum_{k=0}^{T-1}\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ (2a)
+$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
+$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (2b)
+$x_{0}\sim\mathcal{N}(\overline{\mu}_{0},$ $\overline{\Sigma}_{0}),\;\mathbb{E}[x_{T}]=\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$ (2c)
+$P(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad P(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
+where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\Omega$. The distribution $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
 Assumption 1.
-$\rho_{k}(i)>0,\;\forall k\in\mathbb{Z}_{0}^{T},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
+$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
 Assumption 2.
-The covariance of the given MJLS is positive definite at all time steps, i.e. $\Sigma_{k}\succ 0$ $\forall k\in\mathbb{Z}_{0}^{T}$.
-II-A Mean Propagation
-We now focus on MJLS mean dynamics. We know:
-$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
-$=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
+$\Sigma_{k}\succ 0$ $\forall k\in{\mathbb{Z}_{0:T}}$.
+Problem 1 is hard to solve because of stochastic decision variables.
+To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS.
+We consider a mode-dependent affine feedback control policy:
+${u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)\left(x_{k}-\overline{x}_{k}(i)\right)}$ (3)
+Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously.
+Focusing on the mean propagation first, we consider $q_{k}(i)$ below:
+$q_{k+1}$ $(j)=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
 $\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
@@
 $\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (5)
-We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy eq. 4.
-II-B Covariance Propagation
-For covariance propagation, we first consider $\mathcal{X}_{k+1}(j)$:
-$\mathcal{X}_{k+1}(j)=\mathbb{E}\left[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}\right]$ (5)
-$=\sum_{i\in\Omega}\left\{p_{ij}\mathbb{E}\left[\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)\right.\right.$
-$\left.\left.\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)^{\top}\mathds{1}_{r_{k}=i}\right]\right\}$
-Using identities derived in the appendix, we can simplify eq. 5, and substitute that in $\Sigma_{k+1}=\sum_{j\in\Omega}\mathcal{X}_{k+1}(j)-\mu_{k+1}\mu_{k+1}^{\top}$ to get:
+Similarly, covariance propagation is studied below:
+$\Sigma_{k+1}$ $=\left(\sum_{i\in\Omega}\mathbb{E}[(x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k}=i}]\right)-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
+Substituting the identities given in the appendix, we get:
 ${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
@@
 Remark 1.
-Separately analyzing some terms inside the summation in eq. 6:
-$A_{K}(i)\mu_{k}\mu_{k}^{\top}A_{K}^{\top}(i)+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}$
-$+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)$
-$+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}B_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
-$=(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})^{\top}$
-Note these terms are linear in $p_{ij}$ while $\mu_{k+1}\mu_{k+1}^{\top}$ involves the product $p_{ij}p_{ij}$, explaining why they don’t cancel out.
-This demonstrates why the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
+eq. 7 shows that the mean-covariance separation of single-mode dynamics ([10]) does not extend to MJLS.
+We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in{\mathbb{Z}_{0:T-1}}}$ which satisfy eq.
+5.
+We now define our optimal control problem in terms of $\mu_{k}$ and $\Sigma_{k}$, which approximates Problem 1.
+Problem 2.
+Solve:
+$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (8)
+$s.t.\;\forall k\in\mathbb{Z}_{0:T-1},$
+[…]
+We write $J_{k}$ in terms of $\mu_{k}$ and $\Sigma_{k}$ as:
+$J_{k}$ $=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
+Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
+$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (9)
+$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
+The nonlinearity in eq.
+7 make Problem 2 results in a non-convex optimization landscape, while the chance constraints eq.
+2d are not yet represented in closed form.
+We tackle these challenges in the following sections by first proposing a tight convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain tight with a deterministic formulation of chance constraints too.
 III Unconstrained MJLS Covariance Steering
@@
 The unconstrained optimal control problem is given by:
-$\min$ $\;J=\sum_{k=0}^{T-1}J_{k}$ (7)
-$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},$
-$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+G_{k}(r_{k})c_{k}$
-$x_{0}$ $\sim\mathcal{N}(\overline{\mu}_{0},\overline{\Sigma}_{0})$
-$\mathbb{E}[x_{T}]=$ $\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$
-where $J_{k}=\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ with $Q_{k}\succeq 0$ and $R_{k}\succ 0$, $\forall k\in\mathbb{Z}_{0}^{T-1}$. Simplifying $J_{k}$, we get:
-$J_{k}$ $=\sum_{i\in\Omega}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
-$=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
-Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
-$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (8)
-$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
+Problem 3.
+Solve:
+$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (10)
+[…]
+$\;\;\mu_{0}=\overline{\mu}_{0}\;\;\mu_{T}=\mu_{f}$ (11)
 III-B Convex Relaxation
-The nonlinear terms in eq.
-6 result in a non-convex optimization landscape for eq.
-7.
-To solve it in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations.
+To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations.
 We first introduce the following change of variables to convexify eq.
-6:
+7:
 $L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (12)
@@
 Thus, section III-B can be expressed as
-$\Sigma_{k+1}=F_{\tau}(\Xi)$ (11)
+$H_{k}^{\Sigma}=S_{k+1}(j)-F_{\tau}(\Xi)=0$ (14)
 where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 12 as:
@@
 III-C Two-step Convex Optimization Framework
-We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 1 and then uses its solution to solve Problem 2.
-Problem 1.
+We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 4 and then uses its solution to solve Problem 5.
+Problem 4.
 Solve:
@@
 $\mu_{0}=\overline{\mu}_{0},\;\mu_{T}=\mu_{F}$ (19c)
-Problem 2.
+Problem 5.
 Given a mean trajectory $\tau$, solve:
@@
 […]
-$H_{K}^{\Sigma}=\Sigma_{k+1}$ $-F_{\tau}(\Xi)=0$ (17b)
-$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (17c)
-$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (17d)
+$\Sigma_{0}=\overline{\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (20c)
 Assumption 3.
-Slater’s condition is satisfied for Problem 2, i.e. there exists a strictly feasible solution for Problem 2.
+Slater’s condition is satisfied for Problem 5, i.e. there exists a strictly feasible solution for Problem 5.
 Lemma 1 (Lemma 1 in [11]).
@@
 Theorem 1.
-Under Assumptions 1, 2 and 3, the optimal solution to Problem 2 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
+Under Assumptions 1, 2 and 3, the optimal solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
 Proof.
-For Problem 2, we can define the Lagrangian as:
+For Problem 5, we can define the Lagrangian as:
 $\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (21)
@@
 where $C_{k}(i)$ is defined in eq.
-12, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively.
+15, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively.
 $\Lambda_{k}$ is symmetric because of symmetry in eq.
-17b which shares the same symmetry properties with eqs.
-11 and III-B.
-$M_{k}$ and $\lambda_{F}$ are symmetric by definition.
+14 and $M_{k}$ and $\lambda_{F}$ are symmetric by definition.
 The relevant first-order KKT conditions for optimality are:
@@
 Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq.
-19c.
+22c.
 Rearranging eq.
-19a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
+22a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
 Using this and the substitutions defined earlier in eq.
-9, we simplify eq.
-19b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
+12, we simplify eq.
+22b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
 Taking the determinant of both sides, we get the following:
@@
 Remark 2.
-If we attempt to carry out similar relaxations for the mean quadratic terms to relax the covariance propagation equation (6) without assuming the mean trajectory $\tau$ is given, we find that it is not possible to prove that these relaxations are lossless.
-In fact, we get a condition where $M_{k}^{(3)}=-M^{(4)}_{k}$, where $M_{k}^{(3)}$ and $M^{(4)}_{k}$ are Lagrange multipliers for respective relaxations.
-Since they are both positive semi-definite, it implies $M_{k}^{(3)}=M^{(4)}_{k}=0$.
-Thus, we cannot use our existing arguments to prove the relaxation is lossless.
-Numerical simulations also demonstrated this lossy behavior, motivating our two-step sequential mean-covariance optimization approach.
+Similar relaxations cannot be performed for mean quadratic terms in the eq.
+7 because they involve rank-1 matrices which doesn’t allow the use of Schur’s complement to relax the non-convex substitutions as LMIs, motivating our two-step sequential mean-covariance optimization approach.
 IV Chance-Constrained MJLS Covariance Steering
@@
 Previous works [24, 8] have focused on deterministic and convex reformulations of eqs.
-20 and 21, but they usually require Gaussian dynamics.
+23 and 24, but they usually require Gaussian dynamics.
 MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution.
@@
 We now discuss the deterministic formulation of eqs.
-20 and 21.
+23 and 24.
 IV-A Deterministic Formulation
@@
 Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
-Note that while eqs. 22 and 23 are convex in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
+Note that while eqs. 25 and 26 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
 $m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (27)
@@
 We now define the relaxed mean and covariance problems as follows:
-Problem 3.
+Problem 6.
 Given a covariance trajectory $\Xi$, solve:
-$\min_{\mu_{k},\overline{u}_{k}(i),\beta_{m}}J_{\mu}^{c}$ (26a)
+$\min_{\mu_{k},\overline{u}_{k}(i)}J_{\mu}^{c}$ (29a)
+[…]
 […]
@@
 where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq.
-15a.
+18a.
 $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$.
 $\alpha_{m}>0$ is the weight corresponding to the slack variable vector 1-norm $\|\beta_{m}\|_{1}$ in $J_{\mu}^{c}$.
-Problem 4.
+Problem 7.
 Given a mean trajectory $\tau$, solve:
-$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i),\zeta_{m}}J^{c}_{\Sigma}\hskip 56.9055pt$ (27a)
+$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J^{c}_{\Sigma}\hskip 56.9055pt$ (30a)
 […]
-$H_{K}^{\Sigma}=\Sigma_{k+1}-F_{\tau}(\Xi)$ $=0$ (27b)
-$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (27c)
-$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (27d)
-$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (27e)
-$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $\leq\zeta_{2,j},$ (27f)
-$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (27g)
-$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (27h)
+[…]
+$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $-\zeta_{1,j}\leq 0,$ (30c)
+$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $-\zeta_{2,j}\leq 0,$ (30d)
+$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $-\zeta_{3,j}\leq 0$ (30e)
+$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $-\zeta_{4,j}\leq 0$ (30f)
 where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq.
-15b.
+18b.
 $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$.
@@
 Assumption 4.
-Slater’s condition is satisfied in Problem 4 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 27c remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
-We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 4.
+Slater’s condition is satisfied in Problem 7 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 16 remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
+We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 7.
 Theorem 2.
-Under Assumptions 1, 2 and 4, the optimal solution to Problem 4 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
+Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
 Proof.
-We can define the Lagrangian for Problem 4 as follows:
+We can define the Lagrangian for Problem 7 as follows:
 $\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl{[}\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (31)
@@
 15 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs.
-27e, 27f, 27g and 27h respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers.
+30c, 30d, 30e and 30f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers.
 $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1.
@@
 Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq.
-29c.
+32c.
 Rearranging eq.
-29a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
+32a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
 Using this and the substitutions defined in eq.
-9, we simplify eq.
-29b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
-Representing the left hand side as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
+12, we simplify eq.
+32b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
+Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
 $\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
@@
 We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints.
-The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 1) and then solve Problem 4 and Problem 3 iteratively with increasing weights on the slack variables until they reach negligible values.
+The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 4) and then solve Problem 7 and Problem 6 iteratively with increasing weights on the slack variables until they reach negligible values.
 This framework is described in Algorithm 1.
@@
 1:Set tol, initial weights $\{\alpha_{m}\}$, and $\eta>1$;
-2:Initialize mean trajectory $\tau$ by solving Problem 1;
+2:Initialize mean trajectory $\tau$ by solving Problem 4;
 3:while NotConverged do
-4:     Solve Problem 4 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
-5:     Compute new mean trajectory $\tau$ by solving Problem 3 using revised covariance trajectory $\Xi$ from previous step ;
+4:     Solve Problem 7 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
+5:     Compute new mean trajectory $\tau$ by solving Problem 6 using revised covariance trajectory $\Xi$ from previous step ;
 6:     Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
@@
 We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$.
-We set the cost weight matrices $Q_{k}=0$ and $R_{k}=I$ $\forall k\in\mathbb{Z}_{0}^{T-1}$.
 We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM.
@@
 Numerical simulations for the covariance steering framework are presented.
+Future work will analyze convergence behavior of the framework and extensions to semi-markov jump systems.
 Appendix A MJLS identities
@@
 $=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (35)
+Appendix B Discussion on independence of $x_{k}$ and $r_{k}$
+In this section, we provide numerical results showing that the relationship between random variables $x_{k}$ and $r_{k}$ is neglible and they can therefore, be considered independent for all practical purposes.
+We first compare $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$.
+A close agreement between the two would indicate our initial assumption.
+We consider a scalar uncontrolled ($B_{k}(i)=0$) MJLS system without noise or bias ($G_{k}(i)=c_{k}=0$) to simplify the analysis.
+The system parameters are given below:
+$A_{k}(1)=0.8$ $\quad A_{k}(2)=1.2$ $\mathbb{P}=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $\quad\rho_{k}(i)=[0.7;0.3]$
+Figure 3 shows three plots which compare the sample mean $\mu_{k}$ multiplied by $\rho_{k}(i)$ to $q_{k}(i)$ which is calculated analytically using eq.
+4.
+The plots show high agreement between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$ indicating that the dependence between $\mathds{1}_{r_{k}=i}$ and $x_{k}$ is neglible.
+Figure 3: Figure shows comparison between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$, where $\mu_{k}$ is calculated using samples and $q_{k}(i)$ is calculated analytically for both modes.
+The right most plot shows the comparison of sampled and analytical mean $\mu_{k}$, demonstrating that the number of samples considered are sufficient for statistical convergence.
+To further demonstrate this, and also provide numerical evidence to prove our mean and covariance propagation equations are correct, we consider the following time-varying system:
+$A_{1}(k)$ $=\begin{bmatrix}-0.5&1\\ -1&\sin(0.2\pi k)\end{bmatrix}$ $A_{2}(k)$ $=\begin{bmatrix}\cos(0.2\pi k)&0.5\\ 0.5&0.5\end{bmatrix}$ (36)
+$B_{1}(k)$ $=\begin{bmatrix}1\\ 0\end{bmatrix}$ $B_{2}(k)$ $=\begin{bmatrix}0.5\\ 1\end{bmatrix}$
+$G_{1}(k)$ $=\begin{bmatrix}1\\ 1.2\end{bmatrix}$ $G_{2}(k)$ $=\begin{bmatrix}0.2\\ 0.3\sin(0.2\pi k)\end{bmatrix}$
+$T$ $=10$ $\mathbb{P}$ $=\begin{bmatrix}0.6&0.4\\ 0.5&0.5\end{bmatrix}$
+We use a randomized open loop control law sampled uniformly from $[-10,10]$.
+The noise is $w_{k}\sim\mathcal{N}(1,1000)$ and the initial state distribution is given by $x_{0}\sim\mathcal{N}(\begin{bmatrix}20,100\end{bmatrix}^{\top},10I_{2})$.
+Figure 4 plots the comparison between sampled mean and covariance values with theoretical calculations using the mean and covariance propagation equations provided in this work for different number of samples.
+Our simulations demonstrate the numerical accuracy of the propagation equations, while also validating our assumption about $x_{k}$ and $r_{k}$.
+Figure 4: Comparison of sampled mean and covariance values with theoretical values calculated using propagation equations for 10,100 and 1e5 samples.

### v4

@@
 Abstract
-In this paper, we propose a novel convex optimization framework to solve the optimal covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) with chance constraints.
+In this paper, we solve the chance-constrained covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) using a convex optimization framework.
 We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case.
 To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially.
-Further, we use Gaussian approximations to incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem.
+Further, we incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem for MJLS.
 Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions.
-Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under Gaussian noise and Markovian jump dynamics.
+Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under additive noise, bias, and Markovian jump dynamics.
 I Introduction
-Covariance steering is an important problem of designing a control policy that steers a system from an initial distribution to a target distribution while minimizing a cost function.
-It was first proposed in [1] as an infinite horizon problem and has attracted significant interest recently[2, 3, 4, 5].
+Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function.
+It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
 The covariance steering formulation was previously applied to vehicle path planning [6] and spacecraft trajectory optimization[7, 8, 9].
@@
 Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored.
-[13] studies covariance steering for a subclass of hybrid systems but don’t extend to Markov jump linear systems (MJLS).
-MJLS is a type of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain.
-They have been extensively studied in the literature[14] with applications in fault-tolerant control [15], networks [16], and macroeconomic models [17].
-Constrained control of MJLS has been an active area of research, and has benefited from the enormous progress in constrained control of single mode systems [bemporad_explicit_2002, 18] e.g.
+[13] studies covariance steering for a class of hybrid systems that involve state-triggered switching.
+However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS).
+MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain.
+They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16].
+[17] is an excellent reference on the classical results for discrete-time MJLS.
+Constrained control of MJLS has been an active area of research, e.g.
 with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20].
 Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons.
-Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints in a computationally tractable manner.
+Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner.
 To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
 In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints.
-This is motivated by the need for probabilistic path planning approaches which face abrupt changes in dynamics, such as spacecrafts[23] and robots[21] with sudden actuator failure.
+This is motivated by the need for probabilistic motion planning approaches which face abrupt changes in dynamics, such as spacecrafts[22] and robots[20] with sudden actuator failure.
 The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
 Notation
-[…]
+$\mathbb{R}$ and $\mathbb{Z}_{a:b}$ denote the set of real numbers and integers from $a$ to $b$, respectively.
+$I$ denotes the identity matrix of appropriate size.
+$\mathds{1}_{\mathcal{\omega}}$ denotes the indicator function.
+$\lambda_{\max}(\cdot)$, $\mathrm{det}(\cdot)$, and $\mathrm{Tr(\cdot)}$ denote the maximum eigenvalue, determinant, and trace of a matrix respectively.
+$\|\cdot\|$ denotes the $l_{2}$ norm for a vector, while $\|\cdot\|_{p}$ denotes its $l_{p}$ norm.
 II Problem Statement
-We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
+We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a wide-sense stationary white noise $w_{k}\in\mathbb{R}^{n_{w}}$ and a deterministic bias $c_{k}$, given by:
 $x_{k+1}=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (1)
 where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively.
-$r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in $\Omega=\{1\ldots N\}$ with transition probability matrix $\mathbb{P}$ whose ($i$,$j$)th entry is $p_{ij}$.
-We define the following variables to study MJLS dynamics:
+$r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in the set $\mathcal{S}\triangleq\{1\ldots N\}$ with transition probability matrix $\mathcal{P}$ whose (ii,$j$)th entry is $p_{ij}$.
+For the white noise $w_{k}$, we have $\mathbb{E}[w_{k}]=0$, $\mathbb{E}[w_{k}w^{\top}_{k}]=I_{n_{w}}$, and $\mathbb{E}[w_{j}w^{\top}_{k}]=0\;\forall i,j,k,j\neq k$ by definition.
+To study the MJLS dynamics, we focus on the mode probability $\rho_{k}(i)$, and the MJLS state mean $\mu_{k}$ and covariance $\Sigma_{k}$.
+In MJLS literature, it has typically been easier to work with partial expectations such as $\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]$ instead of $\mathbb{E}[x_{k}]$ and we define such variables for the mean and covariance below:
 $\rho_{k}(i)\triangleq\mathbb{E}\left[\mathds{1}_{\{r_{k}=i\}}\right],$ $q_{k}(i)\triangleq\mathbb{E}\left[x_{k}\mathds{1}_{r_{k}=i}\right],$
-$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
+$\overline{x}_{k}(i)\triangleq\mathbb{E}[x_{k}|r_{k}=i],$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
+$\mu_{k}\triangleq\mathbb{E}[x_{k}],$ $\hskip-34.1433ptS_{k}(i)\triangleq\mathbb{E}[(x-\overline{x}_{k}(i))(x-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]$
 The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$.
-For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
-We approximate $q_{k}(i)$ by $\rho_{k}(i)\mu_{k}$, which was found to be a reasonable approximation numerically.
-Similarly, we approximate $\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]$ by $\rho_{k}(i)\Sigma_{k}$.
-The appendix section elaborates on this approximation and the consequent justification.
-We now define the MJLS covariance steering optimal control problem below:
+For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathcal{P}^{k}$.
+Our framework can be extended trivially to time-inhomogeneous MJLS with $\rho_{k}=\rho_{0}\mathcal{P}_{0}\mathcal{P}_{1}\ldots\mathcal{P}_{k-1}$, but we only consider the time-homogeneous case for notational simplicity.
+The MJLS covariance steering optimal control problem is then defined as:
 Problem 1.
@@
 $\min_{x_{k},u_{k}(i)}\;J$ $=\sum_{k=0}^{T-1}\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ (2a)
-$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
+$\mathrm{s.t}$ $.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
 $x_{k+1}=A_{k}$ $(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (2b)
-$x_{0}\sim\mathcal{N}(\overline{\mu}_{0},$ $\overline{\Sigma}_{0}),\;\mathbb{E}[x_{T}]=\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$ (2c)
-$P(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad P(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
-where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\Omega$. The distribution $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
+$\mathbb{E}[x_{0}]=\mu_{0}$ $,\;\mathrm{cov}(x_{0})={\Sigma}_{0},\;\mathbb{E}[x_{T}]=\mu_{f},\;\mathrm{cov}(x_{T})\preceq\Sigma_{f}$ (2c)
+$\mathbb{P}(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad\mathbb{P}(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
+where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\mathcal{S}$.
+The distribution for $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
+${\mu}_{0},\;{\Sigma}_{0}$ are the initial mean and covariance, while $\mu_{f},\Sigma_{f}$ are the final time mean and covariance constraint with $\mathrm{cov}(x)$ denoting the covariance of $x$.
+$\mathcal{X}$ and $\mathcal{U}$ indicate the feasible sets which must be satisfied with probability $1-\epsilon_{x}$ and $1-\epsilon_{u}$ for state and control respectively.
 Assumption 1.
-$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
-Assumption 2.
-$\Sigma_{k}\succ 0$ $\forall k\in{\mathbb{Z}_{0:T}}$.
-Problem 1 is hard to solve because of stochastic decision variables.
+$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\mathcal{S}$, i.e. the probability of any mode being active is non-zero at all times.
+Problem 1 is hard to solve as the decision variables are random variables.
 To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS.
+However, instead of dealing with $\mu_{k}$ and $\Sigma_{k}$, the MJLS dynamics motivate us to instead work with $q_{k}(i),\overline{x}_{k}(i)$ and $S_{k}(i)$, which are related to $\mu_{k}$ and $\Sigma_{k}$ as:
+$\mu_{k}$ $=\sum_{i\in\mathcal{S}}q_{k}(i),\;\;\overline{x}_{k}(i)=\frac{q_{k}(i)}{\rho_{k}(i)}$ (3a)
+$\Sigma_{k}$ $=\sum_{i\in\mathcal{S}}\left(S_{k}(i)+\rho_{k}(i)(\overline{x}_{k}(i)-\mu_{k})(\overline{x}_{k}(i)-\mu_{k})^{\top}\right)$ (3b)
 We consider a mode-dependent affine feedback control policy:
 $u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)\left(x_{k}-\overline{x}_{k}(i)\right)$} (4)
+Assumption 2.
+Under eq. 4, $S_{k}(i)\succ 0,\;\forall k\in{\mathbb{Z}_{0:T},i\in\mathcal{S}}$.
 Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously.
-Focusing on the mean propagation first, we consider $q_{k}(i)$ below:
-$q_{k+1}$ $(j)=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
+We will now use the first and second order moments to solve Problem 1 in a tractable manner.
+Focusing on the mean, $q_{k}(i)$ propagation can be performed as given below:
+$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
+[…]
 $\left.\left.\hskip 85.35826pt+c_{k}+G_{k}(i)w_{k}\right)p_{ij}\mathds{1}_{\{r_{k}=i\}}\right]$
-$=\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (4)
-Using the definition of $\mu_{k}$, eq. 4 can also be written as:
-$\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (5)
-Similarly, covariance propagation is studied below:
-$\Sigma_{k+1}$ $=\left(\sum_{i\in\Omega}\mathbb{E}[(x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k}=i}]\right)-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
+$\hskip-14.22636pt={\sum_{i\in\mathcal{S}}p_{ij}\left({A}_{k}(i)q_{k}(i)+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)+c_{k})\right)}$ (5)
+Similarly, covariance propagation can be studied using $S_{k}(i)$ as follows:
+$S_{k+1}(j)$ $=\mathbb{E}[(x_{k+1}-\overline{x}_{k+1}(j))(x_{k+1}-\overline{x}_{k+1}(j))^{\top}\mathds{1}_{r_{k+1}=j}]$
+$\hskip-14.22636pt=\mathbb{E}[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}]-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (6)
 Substituting the identities given in the appendix, we get:
-${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
-$\left.+A_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K_{k}^{\top}(i))B^{\top}_{k}(i)+A_{k}(i)\mu_{k}c^{\top}_{k}\right.$
-$\left.+B_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K^{\top}_{k}(i))^{\top}A_{k}^{\top}(i)+B_{k}(i)\overline{u}^{\top}_{k}(i)c^{\top}_{k}\right.$
-$\left.+B_{k}(i)(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)\Sigma_{k}K_{k}^{\top}(i))B_{k}^{\top}(i)+c_{k}c_{k}^{\top}\right.$
-$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
-$\left.+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$ (7)
+$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
+$+A_{k}(i)S_{k}(i)K_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)K_{k}(i)S_{k}^{\top}(i)A_{k}^{\top}(i)$
+$+B_{k}(i)(\rho_{k}(i)\overline{u}_{k}(i)\overline{u}_{k}^{\top}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i))B_{k}^{\top}(i)$
+$+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)c_{k}^{\top}+B_{k}(i)\overline{u}_{k}(i)\overline{x}_{k}^{\top}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
+$+c_{k}\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)+c_{k}\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+A_{k}(i)\overline{x}_{k}(i)\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)$
+$+A_{k}(i)\overline{x}_{k}(i)c_{k}^{\top}+A_{k}(i)\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)$
+$\hskip 71.13188pt-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (7)
 Remark 1.
-eq. 7 shows that the mean-covariance separation of single-mode dynamics ([10]) does not extend to MJLS.
-We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in{\mathbb{Z}_{0:T-1}}}$ which satisfy eq.
-5.
-We now define our optimal control problem in terms of $\mu_{k}$ and $\Sigma_{k}$, which approximates Problem 1.
+A careful analysis of eq.
+7 along with eq.
+3b show that the terms involving the MJLS mean information do not cancel out for the covariance propagation, demonstrating that the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
+For a given finite sequence of tuples $\{q_{k}(1)\ldots q_{k}(N)\}_{k\in\mathbb{Z}_{0:T}}$ and $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy eq.
+5, we define a corresponding mean trajectory $\tau$ as a set of two finite sequences of $N+1$ and $N$-tuples $\{\mu_{k},\overline{x}_{k}(1)\ldots\overline{x}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T}}}$ $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T-1}}}$ respectively which satisfy eq.
+3a and eq.
+5 simultaneously.
+We now redefine the cost function $J$ in terms of the newly defined mean and covariance variables. From eq. 2a, we know $J=\sum_{k=0}^{T-1}J_{k}$, where $J_{k}$ is given by
+$J_{k}$ $=\sum_{i\in\mathcal{S}}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
+$=\sum_{i\in\mathcal{S}}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
+Using the identities derived in appendices A and A in the appendix, rewrite $J_{k}$ as:
+$J_{k}$ $=\sum_{i\in\mathcal{S}}\left[\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}\right$.
+$\left.\hskip 14.22636pt+\mathrm{Tr}({S}_{k}(i)Q_{k})+\mathrm{Tr}(K_{k}(i)S_{k}(i)K^{\top}_{k}(i)R_{k})\right]$ (8)
+We now define our optimal control problem in terms of the mean and covariance variables using eq. 4, which approximates Problem 1.
 Problem 2.
 Solve:
-$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (8)
-$s.t.\;\forall k\in\mathbb{Z}_{0:T-1},$
+$\min_{\mu_{k},q_{k}(i),\overline{x}_{k}(i),\overline{u}_{k}(i),K_{k}(i),S_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (9)
+$\mathrm{s.t.}\;\forall k\in\mathbb{Z}_{0:T-1},$
 […]
-We write $J_{k}$ in terms of $\mu_{k}$ and $\Sigma_{k}$ as:
-$J_{k}$ $=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
-Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
-$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (9)
-$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
-The nonlinearity in eq.
-7 make Problem 2 results in a non-convex optimization landscape, while the chance constraints eq.
-2d are not yet represented in closed form.
-We tackle these challenges in the following sections by first proposing a tight convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain tight with a deterministic formulation of chance constraints too.
+Problem 2 is still not a tractable problem to solve because of the nonlinear constraint eq.
+7 and chance constraints eq.
+2d, which are not yet defined as a deterministic constraint in the mean and covariance variables.
+We tackle these challenges in the following sections by first proposing a lossless convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain lossless with the deterministic formulations of chance constraints.
+Our lossless relaxations convert an equality constraint to an inequality constraint, which is then shown to be tight, i.e.
+equivalent to the original equality constraint at optimality.
 III Unconstrained MJLS Covariance Steering
@@
 Problem 3.
-Solve:
-$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (10)
-[…]
-$\;\;\mu_{0}=\overline{\mu}_{0}\;\;\mu_{T}=\mu_{f}$ (11)
+Solve Problem 2 without constraints in eq. 2d.
 III-B Convex Relaxation
-To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations.
-We first introduce the following change of variables to convexify eq.
-7:
-$L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (12)
+The nonlinear terms in eq.
+7 result in a non-convex optimization landscape for Problem 3.
+We first relax the nonlinear terms involving the covariance variables $K_{k}(i)$ and $S_{k}(i)$ in eq.
+7 by introducing the following change of variables:
+$L_{k}(i)$ $=K_{k}(i){{S}_{k}(i)},\;$ $Y_{k}(i)$ $=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$ (10)
 Rearranging eq. 7 using the above-mentioned terms,
-${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
-$\left.+A_{k}(i)L^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)\right.$
-$\left.+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B^{\top}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}\right.$
-$\left.+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)\right.$
-$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)]\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$
-For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a finite sequence of $(2N+1)$-tuples $\{\Sigma_{k},L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy section III-B.
+$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
+$+A_{k}(i)L_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)$
+$+\rho_{k}(i)B_{k}(i)\overline{u}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
+$+\rho_{k}(i)c_{k}(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
+$+\rho_{k}(i)A_{k}(i)\overline{x}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
+$+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (11)
+For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a set of two finite sequences of $N+1$ and $2N$-tuples $\{\Sigma_{k}\}_{k\in\mathbb{Z}_{0:T}}$ and $\{L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy section III-B.
 Thus, section III-B can be expressed as
@@
 where $F_{\tau}(\Xi)$ is a linear operator in $\Xi$ for a given $\tau$. We relax the non-convex definition of $Y_{k}(i)$ in eq. 10 as:
-$C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (15)
+$C_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)-Y_{k}(i)\preceq 0$ (13)
 which can be recast as a LMI using Schur’s complement:
-$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}\succeq 0$ (16)
-We use these new variables in the objective function as follows:
+$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{{S}_{k}(i)}\end{bmatrix}\succeq 0$ (14)
+We note that section III-B still has nonlinear terms involving the mean variables.
+To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization scheme which solves the mean and covariance problem sequentially.
+The objective function can be divided into a mean and covariance part using the new variables as follows:
 $J$ $=J(\tau)+J(\Xi)$ (15)
 where
-$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right]$ (18a)
-$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}({\Sigma}_{k}Q_{k}+Y_{k}(i)R_{k})\right]$ (18b)
+$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\mathcal{S}}\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}$ (16a)
+$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\mathcal{S}}\left[\text{Tr}({S}_{k}(i)Q_{k}+Y_{k}(i)R_{k})\right]$ (16b)
 III-C Two-step Convex Optimization Framework
-We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 4 and then uses its solution to solve Problem 5.
+We now define the two subproblems which comprise our two-step optimization framework. It first solves Problem 4 and then uses its solution to solve Problem 5.
 Problem 4.
 Solve:
-$\min_{\mu_{k},\overline{u}_{k}(i)}J(\tau)$ (19a)
+$\min_{q_{k}(i),\overline{x}_{k}(i),\overline{u}_{k}(i)}J(\tau)$ (17a)
 […]
-$\mu_{0}=\overline{\mu}_{0},\;\mu_{T}=\mu_{F}$ (19c)
+$q_{0}(i)=\rho_{0}(i){\mu}_{0},\;\sum_{i\in\mathcal{S}}q_{T}(i)=\mu_{F}$ (17c)
 Problem 5.
 Given a mean trajectory $\tau$, solve:
-$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J(\Xi)$ (20a)
+$\hskip-85.35826pt\min_{S_{k}(i),L_{k}(i),Y_{k}(i)}J(\Xi)$ (18a)
 […]
-$\Sigma_{0}=\overline{\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (20c)
+$\hskip-85.35826ptS_{0}(i)=\rho_{0}(i){\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (18c)
+where $\Sigma_{T}$ is calculated with $S_{T}(i)$ using eq. 3b
 Assumption 3.
@@
 Theorem 1.
-Under Assumptions 1, 2 and 3, the optimal solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
+Under Assumptions 1, 2 and 3, the solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$, hence, lossless.
 Proof.
 For Problem 5, we can define the Lagrangian as:
-$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (21)
+$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (19)
 $+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
 where $C_{k}(i)$ is defined in eq.
-15, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively.
+13, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively.
 $\Lambda_{k}$ is symmetric because of symmetry in eq.
@@
 The relevant first-order KKT conditions for optimality are:
-$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (22a)
-$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (22b)
+$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)=0$ (20a)
+$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (20b)
 $\mathrm{Tr}($ $M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0$ (20c)
@@
 Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq.
-22c.
+20c.
 Rearranging eq.
-22a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
+20a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)A_{k}^{-1}(i)$.
 Using this and the substitutions defined earlier in eq.
-12, we simplify eq.
-22b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
+10, we simplify eq.
+20b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
 Taking the determinant of both sides, we get the following:
-$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
-which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
+$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
+which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i)S^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
 Remark 2.
@@
 Mathematically, they’re expressed as:
-$P(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (23a)
-$P(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (23b)
+$\mathbb{P}(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (21a)
+$\mathbb{P}(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (21b)
 Likewise, for control, we consider
-$P(p^{\top}_{j}u_{k}(i)+q_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (24a)
-$P(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (24b)
-Previous works [24, 8] have focused on deterministic and convex reformulations of eqs.
-23 and 24, but they usually require Gaussian dynamics.
-MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution.
-To achieve this, we seek to minimize the Kullback-Leibler divergence [25] between the MJLS distribution and the approximating Gaussian.
-It has been shown previously [26] that this is achieved by performing moment matching, i.e.
-setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated.
-We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions.
-We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian.
-Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$.
-We now discuss the deterministic formulation of eqs.
-23 and 24.
+$\mathbb{P}(f^{\top}_{j}u_{k}(i)+g_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (22a)
+$\mathbb{P}(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (22b)
+Previous works [23, 8, 10, 6] have focused on deterministic and convex reformulations of eqs.
+21 and 22, but they usually assume state distribution to be Gaussian.
+MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) Therefore, we utilize one-sided Chebyshev/Cantelli’s inequality[24] and the multivariate Chebyshev inequality[25] to convert the chance constraints to deterministic path constraints.
+Note that while the constraint on the state $x_{k}$ is mode-independent, the control chance constraints are imposed for a given mode $i$.
+The multi-modal nature of the control variable $u_{k}$ and practical applications motivate us to consider mode-dependent control chance constraints as it would be less conservative than mode-independent control chance constraints.
+For mode-conditioned control chance constraints, the mean and covariance is given by: $\mathbb{E}[u_{k}(i)|r_{k}=i]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]=\frac{Y_{k}(i)}{\rho_{k}(i)}$ as shown in the appendix.
 IV-A Deterministic Formulation
-Lemma 2 (Lemma 2 in [8]).
-Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$ and $\epsilon\in(0,0.5)$. Then, $\mathbb{P}[a_{j}^{\top}{\xi}+b_{j}\leq 0,\;\forall j]\geq 1-\epsilon$ is implied by
-$a_{j}^{\top}\bar{{\xi}}+b_{j}+m_{\mathcal{N}}(\epsilon_{j})\sqrt{a_{j}^{\top}P_{\xi}a_{j}}\leq 0,\;\forall j,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (25)
-where $m_{\mathcal{N}}(\epsilon)=Q_{X\sim\mathcal{N}}(1-\epsilon)$ denotes the quantile function of the standard normal distribution at probability $1-\epsilon$.
-Lemma 3 (Lemma 3 in [8]).
-Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$, $\epsilon\in(0,1)$, and $\gamma>0$. Then, $\mathbb{P}[\|{\xi}\|\leq\gamma]\geq 1-\epsilon$ is implied by
-$\|\bar{{\xi}}\|+m_{\chi^{2}}(\epsilon,n_{\xi})\sqrt{\lambda_{\max}(P_{\xi})}\leq\gamma$ (26)
-$m_{\chi^{2}}(\epsilon,n_{\xi})=\sqrt{Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)}$.
-Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
-Note that while eqs. 25 and 26 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
-$m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (27)
-$m^{2}_{\chi^{2}}(\epsilon,n_{\xi}){\lambda_{\max}(P_{\xi})}\leq(\gamma-\|\bar{\xi}\|)^{2}$ (28)
-which are convex in the covariance $P_{\xi}$.
+Lemmas 2 and 3 are useful in the deterministic formulation of half-space and max-norm bound constraints.
+Lemma 2.
+Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then, $\mathbb{P}[a_{j}^{\top}{v}+b_{j}\leq 0,\forall j]\geq 1-\epsilon$ is implied by:
+$a^{\top}_{j}\overline{v}+b_{j}+\sqrt{\frac{1-\epsilon_{j}}{\epsilon_{j}}a^{\top}_{j}Va_{j}}\leq 0,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (23)
+Proof.
+See Theorem 3.1 in [26], and Lemma 1 in [8]. ∎
+Lemma 3.
+Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then $\mathbb{P}[\|{v}\|\leq v_{\max}]\geq 1-\epsilon$ is implied by:
+$\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$ (24)
+Proof.
+Given that $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$, we know $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})$…(A).
+This is described in Figure 1(a).
+Now using the multivariate Chebyshev inequality [25] for $v$, we get: $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\geq t^{2})\leq\frac{n_{v}}{t^{2}}$…(B).
+Put $t=\sqrt{\frac{n_{v}}{\epsilon}}$, and considering the complement of the event in (B), we get $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon})\geq 1-\epsilon$.
+Note that $(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}$ denotes an ellipsoid centered at $\overline{v}$ with shape matrix $V$.
+We first find the maximum value of $\|v\|$ inside this ellipse.
+Let $\mathcal{E}=\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$, then: $\max_{v\in\mathcal{E}}\|v\|\leq\max_{v\in\mathcal{E}}\|v-\overline{v}\|+\|\overline{v}\|=\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|$.
+Now construct a ball $\mathcal{B}_{v}$ centered at origin with radius $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}$.
+Geometrically, this is shown in Figure 1(b) for a $n_{v}=2$.
+Then, $\mathbb{P}(v\in\mathcal{B}_{v})=\mathbb{P}(\|v\|\leq\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|)\geq\mathbb{P}(v\in\mathcal{E})$ since $\mathcal{E}\subseteq\mathcal{B}_{v}$.
+But $\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$.
+Therefore, $\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\right)\geq\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$.
+Combining this with (A), we get $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})\geq 1-\epsilon$.
+$\overline{v}||+\sqrt{\frac{n}{\epsilon}\lambda_{max}(V)}$$v_{\max}$v||v||p(){p($\cdot$)}
+[…]
+Figure 1: Schematics for the proof of Lemma 3.
+Figure (a) shows that $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n}{\epsilon}\lambda_{\max}(V)}\right)$.
+Figure (b) shows that $\mathbb{P}(v\in\mathcal{B}_{v})\geq\mathbb{P}(v\in\mathcal{E})$ where the ball $\mathcal{B}_{v}$ is given by $\{v:\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\}$ and the ellipsoid $\mathcal{E}$ is given by $\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$.
+Remark 3.
+Note that while eqs. 23 and 24 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
+$\frac{1-\epsilon}{\epsilon}a^{\top}_{j}Va_{j}\leq(a_{j}^{\top}\bar{v}+b_{j})^{2}$ (25)
+$\frac{n_{v}}{\epsilon}{\lambda_{\max}(V)}\leq(v_{\max}-\|\bar{v}\|)^{2}$ (26)
+which are convex in the covariance $V$.
 IV-B Iterative Convex Control Synthesis Framework
@@
 Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [27, 28].
-We now define the relaxed mean and covariance problems as follows:
+We now define the relaxed mean and covariance problems in Problem 6 and Problem 7 respectively as follows:
 Problem 6.
@@
 $\min_{\mu_{k},\overline{u}_{k}(i)}J_{\mu}^{c}$ (27a)
+$\mathrm{s.t.}\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\mathcal{S},j$
 […]
-[…]
-$a_{j}^{\top}\mu_{k}$ $+b_{j}+m_{\mathcal{N}}(\delta_{x,j})\sqrt{a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k}$ (29c)
-$p_{j}^{\top}\overline{u}_{k}(i)$ $+q_{j}+m_{\mathcal{N}}(\delta_{u,j}(i))\sqrt{p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (29d)
-$\|\overline{u}_{k}(i)\|+$ $m_{\chi^{2}}(\epsilon_{u}(i),n_{u})\sqrt{\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (29e)
+$a_{j}^{\top}\mu_{k}$ $+b_{j}+\sqrt{\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k},$ (27c)
+$f_{j}^{\top}\overline{u}_{k}(i)$ $+g_{j}+\sqrt{\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (27d)
+$\|\overline{u}_{k}(i)\|+$ $\sqrt{\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (27e)
 where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $J_{\mu}^{c}=J(\tau)+\sum_{m}\alpha_{m}\|\beta_{m}\|_{1}$, with $J(\tau)$ defined in eq.
-18a.
+16a.
 $\beta_{m,k}>0$ is the slack variable for $m$-th constraint at time step $k$ and $\beta_{m}=[\beta_{m,0}\ldots\beta_{m,T-1}]^{\top}$.
@@
 $\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J^{c}_{\Sigma}\hskip 56.9055pt$ (28a)
+$\mathrm{s.t.}\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},\;i\in\mathcal{S},j,\hskip 14.22636pt$
 […]
-[…]
-$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $-\zeta_{1,j}\leq 0,$ (30c)
-$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $-\zeta_{2,j}\leq 0,$ (30d)
-$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $-\zeta_{3,j}\leq 0$ (30e)
-$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $-\zeta_{4,j}\leq 0$ (30f)
+${\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (28c)
+${\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}-(f_{j}^{\top}\mu_{k}+g_{j})^{2}$ $\leq\zeta_{2,j},$ (28d)
+$\frac{n_{x}}{\epsilon_{x}}\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (28e)
+$\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (28f)
 where $\sum_{j}\delta_{x,j}\leq\delta_{x}$, $\sum_{j}\delta_{u,j}(i)\leq\delta_{u}(i)$, $u_{m}=u_{\max}-\|\overline{u}_{k}(i)\|$, and $J_{\Sigma}^{c}=J(\Xi)+\sum_{j}\alpha_{m}\|\zeta_{m}\|_{1}$ with $J(\Xi)$ defined in eq.
-18b.
+16b.
 $\zeta_{m,k}>0$ is the slack variable for $m$-th constraint at time $k$ and $\zeta_{m}=[\zeta_{m,0}\ldots\zeta_{m,T-1}]^{\top}$.
@@
 Theorem 2.
-Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
+Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$.
 Proof.
@@
 $\mathcal{L}(\cdot)=$ $J_{\Sigma}^{c}+\sum_{k=0}^{T-1}\Bigl[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\lambda_{x,n}e_{x,n}+\lambda_{x,h}e_{x,h}$ (29)
-$+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right.$
+$+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right$.
 $\left.+\lambda_{u,h}(i)e_{u,h}(i)\right\}\Bigr]+\mathrm{Tr}(\lambda^{\top}_{F}(\Sigma_{T}-\Sigma_{f}))$
@@
 13 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs.
-30c, 30d, 30e and 30f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers.
-$\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1.
+28c, 28d, 28e and 28f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers.
+$\Lambda_{k}$$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1.
 The relevant first-order KKT conditions for optimality are:
-$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (32a)
-$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})v_{k}(i)v_{k}^{\top}(i)$
-$+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (32b)
+$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i)S_{k}^{-1}(i)$
+$\hskip 199.16928pt=0$ (30a)
+$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)$
+$+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (30b)
 $\mathrm{Tr}(M_{k}^{\top}(i)C_{k}(i))=0,\quad C_{k}\preceq 0,\quad M_{k}\succeq 0,$
@@
 Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq.
-32c.
+30c.
 Rearranging eq.
-32a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$.
-Using this and the substitutions defined in eq.
-12, we simplify eq.
-32b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
+30a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i)S_{k}^{-1}(i)A_{k}^{-1}(i)$.
+Using this and substitutions defined in eq.
+10, we simplify eq.
+30b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$.
 Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
-$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
+$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
 From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$.
-Now we know $\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction.
-Hence, $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\Omega$.
+Now we know $\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction.
+Hence, $C_{k}(i)=L_{k}(i){S}^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\mathcal{S}$.
 ∎
@@
 6:  Store values of slack variables vector $\beta_{m}$ and $\zeta_{m}$;
-7:     if $\max_{m}\|\beta_{m}\|_{\infty}$$>$tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ $>$tol then
+7:  if $\max_{m}\|\beta_{m}\|_{\infty}$>>tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ >>tol then
 8:   increase weights $\{\alpha_{m}\}$ by a factor $\eta$;
@@
 We use the following dynamics:
-$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}0.1&0\\ 0&0.1\end{bmatrix}$
+$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0.5\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$
 $A_{k}(2)$ $=\begin{bmatrix}0.2&0.1\\ -0.5&0.1\end{bmatrix}$ $B_{k}(2)$ $=\begin{bmatrix}0&1\\ -1&2\end{bmatrix}$ $G_{k}(2)$ $=\begin{bmatrix}0.5&0\\ 0&0.5\end{bmatrix}$
-$\mathbb{P}$ $=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
-The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$.
-We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$.
+$\mathbb{P}$ $=\begin{bmatrix}0.8&0.2\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
+The initial and final conditions are set as $\rho_{0}=[0.3,0.7]$, ${\mu}_{0}=[25,40]^{\top}$, ${\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=3I$.
+We set a half-space state chance constraint of $\mathbb{P}(a^{\top}x+b<0)\geq 0.95$, where $a=[0,-1]^{\top}$ and b=10b=-10 and a control norm bound constraint $\mathbb{P}(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=8$.
+We consider a zero-mean gaussian distribution $\mathcal{N}(0,I_{n_{w}})$ for $w_{k}$.
+For sampling $x_{0}$, we also consider a normal distribution with mean $\mu_{0}$ and covariance $\Sigma_{0}$.
 We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM.
-All initial weights are set at $\alpha_{m}=1e2$, with $\eta=5$ and $tol=1e-6$.
-It takes 3.43 seconds and $3$ iterations for Algorithm 1 to converge to the preset tolerance values.
-Figures 1(a) and 1(b) plot 1000 Monte Carlo trajectories for this problem without and with state chance constraints, respectively.
-Both figures use the control norm chance constraint.
+All initial weights are set at $\alpha_{m}=10^{2}$, with $\eta=1.5$ and $tol=10^{-6}$.
+It takes 3.43 seconds and $7$ iterations for Algorithm 1 to converge to the preset tolerance values.
+Figures 2(a) and 2(b) plot 2500 Monte Carlo trajectories for this problem without and with state chance constraints, respectively.
+In both cases, the control norm chance constraint is also imposed.
 The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey.
-The unshaded area indicates the feasible region under chance constraints, enforced using the state half-space constraint.
+The unshaded area indicates the feasible region under chance constraints.
 Figure 3 shows the control norm history for the Monte Carlo simulations.
-None of the samples disobeyed the state chance constraints, while the control norm exceeded $u_{\max}$ in $1.3\%$ of the total sample size, which meets the prescribed risk bound of $5\%$.
+The Monte Carlo simulations disobeyed the chance constraint in $0.01\%$ of total sample size, which meets the prescribed risk bound of $5\%$.
+Figure 4 plots $x_{T}$ for all samples, overlayed with the terminal covariance constraint ellipsoid $\Sigma_{f}$, sample covariance and the predicted covariance $\Sigma_{T}$ from the optimizer.
 These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints.
@@
 (b) with chance constraints
-Figure 1: State trajectories for Monte Carlo simulations with the state chance constraint $P(x_{1}\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 5)\leq 0.95$.
-Figure 2: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{1}\geq-10)\geq 0.95$.
+Figure 2: State trajectories for Monte Carlo simulations with the state chance constraint $\mathbb{P}(x[2]\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 8)\leq 0.95$.
+Figure 3: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{2}\geq-10)\geq 0.95$.
+$u_{pred}(i)$ corresponds to $\|\overline{u}_{k}(i)\|+\sqrt{\tfrac{n_{u}(i)}{\epsilon_{u}}\lambda_{\max}(Y_{k}(i))}$
+Figure 4: Sample points from Monte Carlo simulations at final time step, with sampled covariance covariance shown in blue, predicted covariance from optimizer in shaded green, and terminal covariance constraint shown in black.
+The covariance ellipses are scaled to contain $95\%$ of data of a gaussian distribution with same mean and covariance.
 VI Conclusion
@@
 We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework.
-We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem.
-Numerical simulations for the covariance steering framework are presented.
-Future work will analyze convergence behavior of the framework and extensions to semi-markov jump systems.
+We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and develop an iterative convex optimization framework to solve the chance-constrained covariance steering problem.
+Numerical simulations for the covariance steering framework are presented, which demonstrate the controller’s performance in steering the covariance of a MJLS to within a given terminal covariance bound while obeying chance constraints with additive noise and bias.
 Appendix A MJLS identities
 We show some of the identities which are used in deriving the covariance propagation equation for MJLS.
-$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
-$=\rho_{k}(i)\left[\mu_{k}\overline{u}^{\top}_{k}(i)+{\Sigma}_{k}K^{\top}_{k}(i)\right]$ (33)
-$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
+$\mathbb{E}[u_{k}(i)|r_{k}=i]=\frac{\mathbb{E}[u_{k}(i)\mathds{1}_{k}=i]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{\rho_{k}(i)\overline{u}_{k}(i)}{\rho_{k}(i)}=\overline{u}_{k}(i)$ (31)
+$\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]$
+$=\frac{\mathbb{E}[K_{k}(i)S_{k}(i)K^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{Y_{k}(i)}{\rho_{k}(i)}$ (32)
+$\mathbb{E}[(x_{k}-\overline{x}_{k}(i))\mathds{1}_{r_{k}=i}]=0$ (33)
+$\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(x_{k}-\overline{x}_{k}(i)+\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i)+$
+$\hskip 14.22636pt\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]=S_{k}(i)+0+0+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$
+$=S_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$ (34)
+$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)$
+$\hskip 142.26378pt(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
+$=\mathbb{E}[((x_{k}-\overline{x}_{k}(i))+\overline{x}_{k}(i))(\overline{u}_{k}(i)+$
+$\hskip 85.35826ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
+$=0+S_{k}(i)K^{\top}_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+0$
+$=\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+S_{k}(i)K^{\top}_{k}(i)$ (35)
+$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
 $=\overline{u}_{k}(i)c_{k}^{\top}\mathbb{E}[\mathds{1}_{r_{k}=i}]+0=\overline{u}_{k}(i)c_{k}^{\top}\rho_{k}(i)$ (36)
-$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))(\overline{u}_{k}(i)+$
-$\hskip 142.26378ptK_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
-$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}\mathds{1}_{r_{k}=i}]-$
-$\mu^{\top}_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])K_{k}^{\top}(i)+K_{k}(i)(\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]-\mu_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])\overline{u}_{k}^{\top}(i)$
-$\hskip 42.67912pt+K_{k}(i)\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
-$=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (35)
-Appendix B Discussion on independence of $x_{k}$ and $r_{k}$
-In this section, we provide numerical results showing that the relationship between random variables $x_{k}$ and $r_{k}$ is neglible and they can therefore, be considered independent for all practical purposes.
-We first compare $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$.
-A close agreement between the two would indicate our initial assumption.
-We consider a scalar uncontrolled ($B_{k}(i)=0$) MJLS system without noise or bias ($G_{k}(i)=c_{k}=0$) to simplify the analysis.
-The system parameters are given below:
-$A_{k}(1)=0.8$ $\quad A_{k}(2)=1.2$ $\mathbb{P}=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $\quad\rho_{k}(i)=[0.7;0.3]$
-Figure 3 shows three plots which compare the sample mean $\mu_{k}$ multiplied by $\rho_{k}(i)$ to $q_{k}(i)$ which is calculated analytically using eq.
-4.
-The plots show high agreement between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$ indicating that the dependence between $\mathds{1}_{r_{k}=i}$ and $x_{k}$ is neglible.
-Figure 3: Figure shows comparison between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$, where $\mu_{k}$ is calculated using samples and $q_{k}(i)$ is calculated analytically for both modes.
-The right most plot shows the comparison of sampled and analytical mean $\mu_{k}$, demonstrating that the number of samples considered are sufficient for statistical convergence.
-To further demonstrate this, and also provide numerical evidence to prove our mean and covariance propagation equations are correct, we consider the following time-varying system:
-$A_{1}(k)$ $=\begin{bmatrix}-0.5&1\\ -1&\sin(0.2\pi k)\end{bmatrix}$ $A_{2}(k)$ $=\begin{bmatrix}\cos(0.2\pi k)&0.5\\ 0.5&0.5\end{bmatrix}$ (36)
-$B_{1}(k)$ $=\begin{bmatrix}1\\ 0\end{bmatrix}$ $B_{2}(k)$ $=\begin{bmatrix}0.5\\ 1\end{bmatrix}$
-$G_{1}(k)$ $=\begin{bmatrix}1\\ 1.2\end{bmatrix}$ $G_{2}(k)$ $=\begin{bmatrix}0.2\\ 0.3\sin(0.2\pi k)\end{bmatrix}$
-$T$ $=10$ $\mathbb{P}$ $=\begin{bmatrix}0.6&0.4\\ 0.5&0.5\end{bmatrix}$
-We use a randomized open loop control law sampled uniformly from $[-10,10]$.
-The noise is $w_{k}\sim\mathcal{N}(1,1000)$ and the initial state distribution is given by $x_{0}\sim\mathcal{N}(\begin{bmatrix}20,100\end{bmatrix}^{\top},10I_{2})$.
-Figure 4 plots the comparison between sampled mean and covariance values with theoretical calculations using the mean and covariance propagation equations provided in this work for different number of samples.
-Our simulations demonstrate the numerical accuracy of the propagation equations, while also validating our assumption about $x_{k}$ and $r_{k}$.
-Figure 4: Comparison of sampled mean and covariance values with theoretical values calculated using propagation equations for 10,100 and 1e5 samples.
+$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))(\overline{u}_{k}(i)+$
+$\hskip 113.81102ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
+$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}-\overline{x}^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
+$\hskip 85.35826pt+K_{k}(i)\mathbb{E}[x_{k}-\overline{x}_{k}(i)\mathds{1}_{r_{k}=i}]\overline{u}_{k}^{\top}(i)$
+$\hskip 28.45274pt+K_{k}(i)\mathbb{E}[(x_{k}-\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
+$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+0+0+K_{k}(i)S_{k}(i)K^{\top}_{k}(i)$
+$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i)$ (37)